Mark started his car at 50 mph and half an hour later, Joe started his from the same place at the same direction. If Joe passes Mark 150 miles away, how fast was Joe driving?
Answer by stanbon(75887)

Mark started his car at 50 mph and half an hour later, Joe started his from the same place at the same direction. If Joe passes Mark 150 miles away, how fast was Joe driving?
------------------------
Mark DATA:
distance = 150 miles ; rate = 50 mph ; time = d/r = 150/50 = 3 hrs
---------
Joe DATA:
distance = 150 miles ; rate = r mph ; time = d/r = 150/r hrs
---
Equation:
Mark time - Joe time = (1/2) hr.
3 - (150/r) = 1/2
--------
(150/r) = 5/2
----
5r = 300
r = 60 mph (Joe's rate)
==============
cheers,
Stan H.
================